An aqueous solution of MgCl2 that is 27.5% (mass%) and has a solution density of 1.25 g/mL. MgCl2 has a molar mass of 95.211 g/mol. What is the molarity of this MgCl2 solution?
Answer #1

27.5% by mass means the aqueous solution have 27.5g of MgCl2 and 100-27.5 = 72.5g of water.

If the total mass of solution is 100g.

Density of the solution given is 1.25 g/mL.

Calculate the volume of the solution by using definition of density as follows:-

Density=mass/volume.

1.25g/mL = 100g/volume

Volume= 100/1.25 mL

Volume = 80 mL

Volume of MgCl2 solution is 80mL. i.e. 0.080 L

​​​​​​

mass of MgCl2 present in this solution is 27.5g.

Molar mass given is 95.211 g/mol

Calculate the moles of MgCl2

#moles = mass of MgCl2 present/molar mass of MgCl2

#moles of MgCl2= 27.5/95.211

#moles = 0.289 moles

moles of MgCl2 present in the solution is 0.289 moles

molarity is defined as the number of moles of solute present in 1 litre of the solution.

Molarity of MgCl2 = moles of MgCl2 present/volume of the solution.

Molarity of MgCl2= 0.289/0.080 mol/L

Molarity = 3.610 mol/L

So, the molarity of the solution is 3.610 M.